A student wants to establish a standing wave on a wire 1.8 m long clamped at both ends. If the wave speed is 72 m/s, what is the minimum frequency she should apply to set up standing waves?


20 Hz


12 Hz


24 Hz


15 Hz


18 Hz

Clear selection



1
Expert's answer
2021-05-21T10:43:06-0400

"\lambda= vT=\frac{v}{f};"

"f=\frac{v}{\lambda};"

"\lambda= \frac{2l}{n};"

"f=\frac{v}{2l}*n;"

"\text {as f min } n=1;"

"f=\frac{v}{2l}= \frac{72}{2*1.8} = 20"

"\text{Answer:}20 \ Hz"